# Functors on the category of finite sets revisited

### v2

@@ 11.3. The calculation of $\mathrm{Hom}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})$
$0\rightarrow\Lambda^{\bullet+1}(\overline{P})\rightarrow\mathbb{P}_{\bullet}\rightarrow\overline{P}^{\otimes\bullet}\rightarrow 0$
@@
(11.4) $0\rightarrow\Lambda^{k+1}(\overline{P})\boxtimes\mathsf{sgn}_{k}\rightarrow\Lambda^{k+1}(P^{\mathbf{FA}})\boxtimes\mathsf{sgn}_{k}\rightarrow\Lambda^{k}(\overline{P})\boxtimes\mathsf{sgn}_{k}\rightarrow 0$
@@
Applying the functor $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,\overline{P}^{\otimes*}/\Lambda)$ to the short exact sequences (11.4) yields the short exact sequence of $\mathbbm{k}\mathbf{FB}^{\mathrm{op}}$-bimodules:
$0\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{\bullet}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{\bullet+1}(P^{\mathbf{FA}}),\overline{P}^{\otimes*}/\Lambda)\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{\bullet+1}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)\rightarrow 0,$
@@
By construction of $\mathbb{P}_{\bullet}$, one has an isomorphism $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\Lambda^{*}(\overline{P}))\cong\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\Lambda^{*}(\overline{P})).$ This combined with Lemma 11.7 gives:
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]=[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})]\quad+\quad\sum_{k\geq 1}[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{k}(P^{\mathbf{FA}}),\overline{P}^{\otimes*}/\Lambda)]\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}},$
in which $[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$ indicates that $\mathsf{sgn}_{k-1}$ is considered as a $\mathbbm{k}\mathbf{FB}$-module.
@@
By the above discussion, Lemma 11.7 implies that, on applying the functor $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,\overline{P}^{\otimes*})$ to the short exact sequences (11.4), one obtains the short exact sequence of $\mathbbm{k}\mathbf{FB}$-bimodules
$0\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})\rightarrow\bigoplus_{k\geq 1}\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{k}(\overline{P})\boxtimes\mathsf{sgn}_{k-1},\overline{P}^{\otimes*}/\Lambda)\rightarrow 0$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})]$ $=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}\mathcal{S}(k)\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$
$=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}\sum_{t\in\mathbb{N}}(-1)^{t}[\mathsf{sgn}_{k+t}]_{\mathbf{FB}^{\mathrm{op}}}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}.$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{k}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)]=\Big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\Big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}\quad-\quad\mathcal{S}(k)\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}},$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]=[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\quad+\quad\sum_{k\geq 1}\Big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\Big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}.$
@@
A result equivalent to Theorem 11.9 was first proved in [PV23] by very different methods. An alternative approach is also given in [Pow24a].

### v3

@@ 11.3. The calculation of $\mathrm{Hom}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})$
$0\rightarrow\Lambda^{\bullet+1}(\overline{P})\rightarrow\mathbb{P}_{\bullet}\rightarrow\overline{P}^{\otimes\bullet}\rightarrow 0.$
@@
(11.4) $0\rightarrow\Lambda^{k+1}(\overline{P})\boxtimes\mathsf{sgn}_{k}\rightarrow\Lambda^{k+1}(P^{\mathbf{FA}})\boxtimes\mathsf{sgn}_{k}\rightarrow\Lambda^{k}(\overline{P})\boxtimes\mathsf{sgn}_{k}\rightarrow 0.$
@@
Applying the functor $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,\overline{P}^{\otimes*}/\Lambda)$ to the short exact sequences (11.4) yields the short exact sequence of $\mathbbm{k}\mathbf{FB}^{\mathrm{op}}$-bimodules (writing $\mathrm{Hom}$ for $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}$)
$0\rightarrow\mathrm{Hom}(\Lambda^{\bullet}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)\rightarrow\mathrm{Hom}(\Lambda^{\bullet+1}(P^{\mathbf{FA}}),\overline{P}^{\otimes*}/\Lambda)\rightarrow\mathrm{Hom}(\Lambda^{\bullet+1}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)\rightarrow 0,$
@@
By construction of $\mathbb{P}_{\bullet}$, one has an isomorphism
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\Lambda^{*}(\overline{P}))\cong\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\Lambda^{*}(\overline{P})).$
This combined with Lemma 11.7 gives:
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]$ $=$ $[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})]$
$+\sum_{k\geq 1}[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{k}(P^{\mathbf{FA}}),\overline{P}^{\otimes*}/\Lambda)]\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}},$
in which $\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$ indicates that $\mathsf{sgn}_{k-1}$ is considered as a $\mathbbm{k}\mathbf{FB}$-module.
@@
By the above discussion, Lemma 11.7 implies that, on applying the functor $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,\overline{P}^{\otimes*})$ to the short exact sequences (11.4), one obtains the short exact sequence of $\mathbbm{k}\mathbf{FB}$-bimodules (writing $\mathrm{Hom}$ for $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}$)
$0\rightarrow\mathrm{Hom}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})\rightarrow\mathrm{Hom}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})\rightarrow\bigoplus_{k\geq 1}\mathrm{Hom}(\Lambda^{k}(\overline{P})\boxtimes\mathsf{sgn}_{k-1},\overline{P}^{\otimes*}/\Lambda)\rightarrow 0.$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes\bullet},\overline{P}^{\otimes*})]$ $=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)+\sum_{k\geq 1}\mathcal{S}(k)\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$
$=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)+\sum_{k\geq 1}\sum_{t\in\mathbb{N}}(-1)^{t}[\mathsf{sgn}_{k+t}]_{\mathbf{FB}^{\mathrm{op}}}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}.$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{k}(\overline{P}),\overline{P}^{\otimes*}/\Lambda)]$ $=$ $\big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}$
$\ -\ \mathcal{S}(k)\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}},$
@@
$[\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbb{P}_{\bullet},\overline{P}^{\otimes*})]$ $=$ $[\mathbbm{k}{\mathbf{FS}}]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)$
$+\ \sum_{k\geq 1}\Big{(}[\mathsf{sgn}_{k}\otimes_{\mathfrak{S}_{k}}\mathbbm{k}{\mathbf{FS}}(-,\mathbf{k})]\odot_{\mathbf{FB}^{\mathrm{op}}}\mathcal{S}(0)\Big{)}\boxtimes[\mathsf{sgn}_{k-1}]_{\mathbf{FB}}.$
@@
A result equivalent to Theorem 11.9 was first proved in [7] by very different methods. An alternative approach is also given in [5].